# Combining Neural Fields and Deformation Models for Non-Rigid 3D Motion Reconstruction from Partial Data

### v1

@@ 4.2 Animal Motion Sequences
Table 5 shows quantitative comparative results on the DT4D-A test set which is divided into two folds containing motions and individuals unseen during training, respectively. Our method outperforms all competing methods on completion. This demonstrates that our near-isometric deformation assumption allows to generalise to different shape classes.
Fold Unsup. Input Method IoU $\uparrow$ CD $\downarrow$ Corr $\downarrow$
Back Proj. point cloud LPDC [50] 53.24% 0.03961 0.04452
Mot2VecSets [12] 73.84% 0.01790 0.04221
Unseen Motion OFlow [42] 67.29% 0.02643 0.03812
CaDeX [29] 76.57% 0.01735 0.02970
Mono. Depth TSDF Ours 76.73% 0.00990 0.035641
Back Proj. point cloud LPDC [50] 47.31% 0.04710 0.04672
Mot2VecSets [12] 66.45% 0.01971 0.04600
Unseen Individual OFlow [42] 57.13% 0.03994 0.04525
CaDeX [29] 64.87% 0.02704 0.03558
Mono. Depth TSDF Ours 66.32% 0.01478 0.04850
Table 5: Quantitative comparisons of 4D Shape Reconstruction from monocular depth sequences on DT4D-A [32] (all methods are retrained). Best overall in bold, best amongst unsupervised methods underlined.
@@ 4.3 Ablation Studies
Figure 4: Geometry aligned features $(\mathcal{F}_{i})_{i\in\{1,...,N\}}$ interpolated on the reconstructed surface and reduced using t-SNE [51] to $3$ channels and visualised as colors. The near-isometric deformation constraint enriches $\mathcal{F}_{i}$ with correspondences.
We assess the benefit of the two main components of our method. First, the fusion mechanism (Sec. 3.1.2) where we ablate the feature-fusion in latent space. Second, the inter-frame deformation constraint (Sec. 3.2) where the network is restricted to the feature-fusion based completion module.
Fusion Mechanism Unseen Motion Unseen Individual
@@
91.73% 0.00299 0.01382 90.35% 0.00295 0.01293
90.78% 0.00323 0.01342 90.30% 0.00290 0.01245
Table 6: Ablation result of feature-fusion on D-FAUST [8]. Best in bold.
Benefit of the fusion mechanism Tab. 6 shows that linking observations in feature space allows to improve the temporal coherence of the reconstructions.
Benefit of the deformation model In addition to providing motion information, which is critical for downstream applications, the deformation constraint allows to improve the consistency of the reconstructions as shown on the hands in Fig 4. The left part of Fig 4 visualises the learnt geometry aligned features $(\mathcal{F}_{i})_{i\in\{1,...,N\}}$ as colors after using TSNE [51] to reduce them to 3 channels. When trained with the deformation constraint, these features encompass a temporal dimension: corresponding shape parts have the same color across time.
@@ 5 Conclusions
We present a novel representation of moving 3D shapes that combines neural distance fields with a mesh deformation model implementing a near-isometric deformation assumption. We experiment on 3D motion reconstruction from monocular depth videos and demonstrate that our representation allows for high fidelity reconstructions and a precise tracking. Our approach can generalise to different shape classes and displays impressive cross-dataset generalisation going beyond results reported in prior works.
When dealing with motions that significantly deviate from the ones seen during training, our approach can provide an inaccurate tracking. Since the tracking strategy is unsupervised, a test-time optimisation strategy can be used to deal with out of distribution motions, which we leave to future works.
@@ 7 Comparison to NPMs
We compare against model-based method Neural Parametric Models (NPMs) [43] on 3D motion reconstruction from monocular depth videos. NPMs learn disentangled pose and shape spaces from a dataset to which partial observations of a moving shape can be fitted during inference through test-time optimisation. We use the pose and shape spaces pre-trained on human shapes from different datasets [32, 37, 38] and provided by the authors. Our data-driven method trains on DFAUST [8]. We tested on the 4DHumanOutfit [5] dataset which consists of sequences of clothed human motions captured using a multi-camera platform. We use the multi-view mesh reconstructions as ground truth for the completion and SMPL [35] fittings as ground truth for the tracking. We synthetically generate monocular depth videos from the mesh sequences and use TSDF volumes as input for both our method and NPMs. Tab. 7 presents the quantitative comparative results. Our method outperforms NPMs in both completion and tracking. Fig. 5 shows a qualitative comparison on two examples. Colors are defined on the first frame and transferred using predicted tracking. Our method achieves higher fidelity completions in both cases. Further, NPMs fail to infer the correct pose in the presence of loose clothing: the legs are crossed in the first example. This shows that model-based strategies struggle with examples that deviate from the shape-pose space hypothesis they consider.
Unsup. Input Method IoU $\uparrow$ CD $\downarrow$ Corr $\downarrow$
Mono. Depth TSDF NPMs [43] 69.84% 0.01192 0.02547
Ours 83.14% 0.00584 0.02410
@@ 8 Ablation of the Deformation Constraint
Tab. 8 shows quantitative results of the ablation of the deformation constraint where the model is restricted to the Feature-Fusion Based Completion Module. The full model is, overall, on par in terms of reconstruction and completion quality while being augmented with crucial motion information. The deformation constraint also promotes the consistency of the reconstructions as shown in Fig. 4 in the main paper.
@@ 9.1.2 Sampling Strategy for SDF Losses
We use ground truth SDF samples generated from the mesh sequences in Eq. 5 and Eq. 6 of the main paper to supervise the completion task. For each frame $i$, we sample a total of $|S|$ points: $30\%$ are sampled uniformly in our bounding box $B$, while $70\%$ are sampled within a distance of $0.05$ to the mesh surface. In our experiments we fix $|S|=50k$. The bounding box’s extents are fixed to $(-0.5,-0.5,-0.5)$ and $(0.5,0.5,0.5)$.
@@ 9.2.1 Association Losses
$\footnotesize\begin{split}l^{cycle}_{2}&=\frac{1}{N-1}(\sum\limits_{i=1}^{N-1}{\|\Pi_{i\rightarrow i+1}(\Pi_{i+1\rightarrow i}C^{i})-C^{i}\|^{2}_{2}}\\ &+\sum\limits_{i=2}^{N}{\|\Pi_{i\rightarrow i-1}(\Pi_{i-1\rightarrow i}C^{i})-C^{i}\|^{2}_{2}}),\end{split}$ (16)
$\footnotesize\begin{split}l^{cycle}_{3}&=\frac{1}{N-2}(\sum\limits_{i=1}^{N-2}{\|\Pi_{i\rightarrow i+2}(\Pi_{i+2\rightarrow i}C^{i})-C^{i}\|^{2}_{2}}\\ &+\sum\limits_{i=3}^{N}{\|\Pi_{i\rightarrow i-2}(\Pi_{i-2\rightarrow i}C^{i})-C^{i}\|^{2}_{2}}),\end{split}$ (17)
@@
where $\Pi_{i\rightarrow i}$ is a self association matrix computed similarly to Eq. 9 and Eq. 10 in the main paper. We also compute associations on the coarse geometry aligned features $(\mathcal{F}^{c}_{i})_{i\in\{1,...,N\}}$ and optimise for these losses that we denote $l^{cycle}_{coarse}$ and $l^{rec}_{coarse}$. We promote these properties for coarse level features; in turn they will be inherited by finer level features i.e. by $(\mathcal{F}_{i})_{i\in\{1,...,N\}}$. The loss $l^{associate}$ in Eq. 11 of the main paper, integrating the coarse level association losses is implemented as follows:
@@ 9.2.2 Deformation Model
We use a patch-based mesh deformation model [11] to model inter-frame evolution. It decomposes a non-rigid deformation into patch-wise rigid deformations blended at the vertex level. Each mesh $\mathcal{M}$ is decomposed into non-overlapping surface patches $(P_{k})_{1\leq k\leq L}$ with their centers $\mathcal{C}=(c_{k}\in\mathbb{R}^{3})_{1\leq k\leq L}$ where $L$ is the number of patches. Each patch $P_{k}$ defines its rigid deformation i.e.  a rotation $R_{k}\in\mathbb{R}^{3\times 3}$ and a translation $u_{k}\in\mathbb{R}^{3}$ as well as a blending function $\alpha_{k}(v)$ that depends on the euclidean distance of $v$ to $c_{k}$. The deformation model optimises for a rigidity constraint that implements a near isometric-deformation assumption by promoting consistent deformations between every patch $P_{k}$ and its neighbours $\mathcal{N}(P_{k})$. The rigidity constraint we use in Eq. 13 in the main paper is implemented as follows:
$l_{rig}(\mathcal{M})=\sum_{(P_{k})_{1\leq k\leq L}}{\sum_{P_{j}\in\mathcal{N}(P_{k})}{\sum_{v\in P_{k}\bigcup P_{j}}{E^{kj}_{v}}}}\mbox{ with,}$, (21)
@@ 9.2.3 Deformation Decoder
To learn this deformation, we employ the deformation decoder introduced for static complete 3D shapes in [39]. It consists of a hierarchical graph convolutional network acting on the patch neighborhoods followed by an MLP. We use three patch levels in the hierarchical graph convolutional network: $20$, $50$ and $400$. It outputs the deformation model’s parameters i.e. $(R_{k}\in\mathbb{R}^{6})_{1\leq k\leq L}$ and new center positions $(u_{k}\in\mathbb{R}^{3})_{1\leq k\leq L}$ for every patch of of $\mathcal{M}_{i}$. To output the deformation parameters induced by the association matrix $\Pi_{i\rightarrow j}$ i.e. the one that deforms $\mathcal{M}_{i}$ into $\mathcal{M}_{j}$, it takes as input patch centers $C^{i}\in\mathbb{R}^{L\times 3}$, patch-wise features $\mathcal{F}^{patch}_{i}\in\mathbb{R}^{L\times C}$, target centers $\Pi_{i\rightarrow j}\mathcal{C}^{j}\in\mathbb{R}^{L\times 3}$, and $\Pi_{i\rightarrow j}\mathcal{F}^{patch}_{j}\in\mathbb{R}^{L\times C}$. Applying the output deformations leads to the deformed shape $\mathcal{M}_{i\rightarrow j}$. The $6D$ representation of rotation [61] is used.
Fig. 8 shows an example of reconstructions and inter-frame deformations computed by the deformation decoder in the case of a fast motion. The top row shows the reconstructions $\mathcal{M}_{i}$, the second row shows the deformation graphs of the patch-wise deformation model, and third and forth rows show the deformed shapes $\mathcal{M}_{i\rightarrow i+1}$ and $\mathcal{M}_{i\rightarrow i-1}$ respectively. Using the inter-frame deformations $\mathcal{M}_{i\rightarrow i+1}$, we extract the tracking using a nearest neighbour search as shown in the bottom row.
@@ 9.3 Training Details
To allow for more stable learning, our network optimises for $l^{network}$ in gradual steps. First, the network only optimises for $l^{coarse}$ to obtain coarse reconstructions. After $N_{1}$ epochs, assuming that we have roughly located the coarse surface, the refinement step is activated and the network optimises for both $l^{coarse}$ and $l^{fine}$ i.e. for $l^{fusion}$. After $N_{2}$ epochs, given that we have converged to good initial refined surfaces, the association search is activated; the network optimises for both $l^{fusion}$ and $l^{associate}$. Finally, after $N_{3}$ epochs, the deformation search i.e. $l^{deform}$ is activated and the network optimises for $l^{network}$ until convergence. Tab. 9 and Tab. 10 detail this in terms of loss weights for the model trained on D-FAUST [8] and the one trained on DT4D-A [32] respectively.
@@
We train our network with the Adam [27] optimizer and use gradient clipping. We use a learning rate of $10^{-3}$ during the first training epoch, $5\times 10^{-4}$ between the $2^{nd}$ and the $60^{th}$ epoch, $2.5\times 10^{-4}$ between the $60^{th}$ epoch and the $100^{th}$ epoch and $1.25\times 10^{-4}$ after the $100^{th}$ epoch.

### v2

@@ 4.2 Animal Motion Sequences
Tab. 5 gives the quantitative results on the DT4D-A test set which is split into motions and individuals unseen during training. Our method outperforms all competing methods on completion showing that the near-isometric deformation assumption allows to generalise to different shape classes.
Fold Unsup. Method IoU $\uparrow$ CD $\downarrow$ Corr $\downarrow$
LPDC [52] 53.24% 0.3961 0.4452
Mot2VecSets [12] 73.84% 0.1790 0.4221
Unseen
Motion
OFlow [44] 67.29% 0.2643 0.3812
CaDeX [30] 76.57% 0.1735 0.2970
Ours 76.73% 0.0990 0.3564
LPDC [52] 47.31% 0.4710 0.4672
Mot2VecSets [12] 66.45% 0.1971 0.4600
Unseen
Individual
OFlow [44] 57.13% 0.3994 0.4525
CaDeX [30] 64.87% 0.2704 0.3558
Ours 66.32% 0.1478 0.4850
Table 5: Quantitative 4D shape reconstruction results from monocular depth on DT4D-A [33] (all methods are retrained). Best overall in bold, best amongst unsupervised methods underlined.
@@ 4.3 Ablation Studies
Figure 5: Geometry aligned features $(\mathcal{F}_{i})_{i\in\{1,...,N\}}$ interpolated on the reconstructed surface and reduced using t-SNE [53] to 33 channels and visualised as colors.
Fusion
Mechanism
Unseen Motion Unseen Individual
@@
91.73% 0.0299 0.1382 90.35% 0.0295 0.1293
90.78% 0.0323 0.1342 90.30% 0.0290 0.1245
Table 6: Ablation of feature-fusion on D-FAUST [8]. Best in bold.
Benefit of the fusion mechanism Tab. 6 shows that linking observations in feature space allows to improve the temporal coherence of the reconstructions. Benefit of the deformation model In addition to providing critical motion information, the deformation constraint improves the reconstructions’ consistency as on the hands in Fig 5. Fig 5 (left) shows the learnt geometry aligned features $(\mathcal{F}_{i})_{i\in\{1,...,N\}}$ as colors . When trained with the deformation constraint, these features encompass a temporal dimension: corresponding shape parts have the same color across time. Quantitative results are in Sec. 8.
@@ 5 Conclusions
We propose a novel representation of moving 3D shapes that combines neural distance fields with a near-isometric mesh deformation model. Our representation allows for high fidelity reconstructions and a precise tracking, as shown in experiments on 4D reconstruction from monocular depth videos. Our approach can generalize to different shape classes and displays impressive cross-dataset generalisation going beyond results reported in prior works.
When dealing with motions that significantly deviate from the ones seen during training, our approach can provide an inaccurate tracking. Since the tracking strategy is unsupervised, test-time optimisation can be used to deal with out-of-distribution motions.
@@ 7 Comparison to NPMs
We compare against model-based method Neural Parametric Models (NPMs) [45] on 3D motion reconstruction from monocular depth videos. NPMs learn disentangled pose and shape spaces from a dataset to which partial observations of a moving shape can be fitted during inference through test-time optimisation. We use the pose and shape spaces pre-trained on human shapes from different datasets [33, 38, 39] and provided by the authors. Our data-driven method trains on D-FAUST [8]. We tested on the 4DHumanOutfit [5] dataset which consists of sequences of clothed human motions captured using a multi-camera platform. We use the multi-view mesh reconstructions as ground truth for the completion and SMPL [36] fittings as ground truth for the tracking. We synthetically generate monocular depth videos from the mesh sequences and use TSDF volumes as input for both our method and NPMs. Tab. 7 presents the quantitative comparative results. Our method outperforms NPMs in both completion and tracking. Fig. 6 shows a qualitative comparison on two examples. Colors are defined on the first frame and transferred using predicted tracking. Our method achieves higher fidelity completions in both cases. Further, NPMs fail to infer the correct pose in the presence of loose clothing: the legs are crossed in the first example. This shows that model-based strategies struggle with examples that deviate from the shape-pose space hypothesis they consider.
Unsup. Method IoU $\uparrow$ CD $\downarrow$ Corr $\downarrow$
NPMs [45] 69.84% 0.1192 0.2547
Ours 83.14% 0.0584 0.2410
@@ 8 Ablation of the Deformation Constraint
We assess the benefit of the two main components of our method. First, the fusion mechanism (Sec. 3.1.2) where we ablate the feature-fusion in latent space. Second, the inter-frame deformation constraint (Sec. 3.2) where the network is restricted to the feature-fusion based completion module. The added benefit of the fusion mechanism is shown in Tab. 6. Tab. 8 shows quantitative results of the ablation of the deformation constraint. The full model is, overall, on par in terms of reconstruction and completion quality while being augmented with crucial motion information. The deformation constraint also promotes the consistency of the reconstructions as shown in Fig. 5.
@@ 9.1.2 Sampling Strategy for SDF Losses
We use ground truth SDF samples generated from the mesh sequences in Eq. 5 and Eq. 6 to supervise the completion task. For each frame $i$, we sample a total of $|S|$ points: $30\%$ are sampled uniformly in the bounding box $B$, while $70\%$ are sampled within a distance of $0.05$ to the mesh surface. In our experiments we fix $|S|=50k$. The bounding box’s extents are fixed to (0.5,0.5,0.5)(-0.5,-0.5,-0.5) and $(0.5,0.5,0.5)$.
@@ 9.2.1 Association Losses
$\begin{split}l^{cycle}_{2}&=\frac{1}{N-1}(\sum\limits_{i=1}^{N-1}{\|\Pi_{i\rightarrow i+1}(\Pi_{i+1\rightarrow i}C^{i})-C^{i}\|^{2}_{2}}\\ &+\sum\limits_{i=2}^{N}{\|\Pi_{i\rightarrow i-1}(\Pi_{i-1\rightarrow i}C^{i})-C^{i}\|^{2}_{2}}),\end{split}$ (16)
$\begin{split}l^{cycle}_{3}&=\frac{1}{N-2}(\sum\limits_{i=1}^{N-2}{\|\Pi_{i\rightarrow i+2}(\Pi_{i+2\rightarrow i}C^{i})-C^{i}\|^{2}_{2}}\\ &+\sum\limits_{i=3}^{N}{\|\Pi_{i\rightarrow i-2}(\Pi_{i-2\rightarrow i}C^{i})-C^{i}\|^{2}_{2}}),\end{split}$ (17)
@@
where $\Pi_{i\rightarrow i}$ is a self association matrix computed similarly to Eq. 9 and Eq. 10. We also compute associations on the coarse geometry aligned features $(\mathcal{F}^{c}_{i})_{i\in\{1,...,N\}}$ and optimise for these losses that we denote $l^{cycle}_{coarse}$ and $l^{rec}_{coarse}$. We promote these properties for coarse level features; in turn they will be inherited by finer level features i.e. by $(\mathcal{F}_{i})_{i\in\{1,...,N\}}$. The loss $l^{associate}$ in Eq. 11, integrating the coarse level association losses is implemented as follows:
@@ 9.2.2 Deformation Model
We use a patch-based mesh deformation model [11] to model inter-frame evolution. It decomposes a non-rigid deformation into patch-wise rigid deformations blended at the vertex level. Each mesh $\mathcal{M}$ is decomposed into non-overlapping surface patches $(P_{k})_{1\leq k\leq L}$ with their centers $\mathcal{C}=(c_{k}\in\mathbb{R}^{3})_{1\leq k\leq L}$ where $L$ is the number of patches. Each patch $P_{k}$ defines its rigid deformation i.e.  a rotation $R_{k}\in\mathbb{R}^{3\times 3}$ and a translation $u_{k}\in\mathbb{R}^{3}$ as well as a blending function $\alpha_{k}(v)$ that depends on the euclidean distance of $v$ to $c_{k}$. The deformation model optimises for a rigidity constraint that implements a near isometric-deformation assumption by promoting consistent deformations between every patch $P_{k}$ and its neighbours $\mathcal{N}(P_{k})$. The rigidity constraint we use in Eq. 13 is implemented as follows:
$l_{rig}(\mathcal{M})=\sum_{(P_{k})_{1\leq k\leq L}}{\sum_{P_{j}\in\mathcal{N}(P_{k})}{\sum_{v\in P_{k}\bigcup P_{j}}{E^{kj}_{v}}}}\mbox{ with,}$ (21)
@@ 9.2.3 Deformation Decoder
To learn this deformation, we employ the deformation decoder introduced for static complete 3D shapes in [40]. It consists of a hierarchical graph convolutional network acting on the patch neighborhoods followed by an MLP. We use three patch levels in the hierarchical graph convolutional network: $20$, $50$ and $400$. It outputs the deformation model’s parameters i.e. $(R_{k}\in\mathbb{R}^{6})_{1\leq k\leq L}$ and new center positions $(u_{k}\in\mathbb{R}^{3})_{1\leq k\leq L}$ for every patch of $\mathcal{M}_{i}$. To output the deformation parameters induced by the association matrix $\Pi_{i\rightarrow j}$ i.e. the one that deforms $\mathcal{M}_{i}$ into $\mathcal{M}_{j}$, it takes as input patch centers $C^{i}\in\mathbb{R}^{L\times 3}$, patch-wise features $\mathcal{F}^{patch}_{i}\in\mathbb{R}^{L\times C}$, target centers $\Pi_{i\rightarrow j}\mathcal{C}^{j}\in\mathbb{R}^{L\times 3}$, and $\Pi_{i\rightarrow j}\mathcal{F}^{patch}_{j}\in\mathbb{R}^{L\times C}$. Applying the output deformations leads to the deformed shape $\mathcal{M}_{i\rightarrow j}$. The $6D$ representation of rotations [64] is used.
Fig. 9 shows an example of reconstructions and inter-frame deformations computed by the deformation decoder in the case of a fast motion. The top row shows the reconstructions $\mathcal{M}_{i}$, the second row shows the deformation graphs of the patch-wise deformation model, and third and forth rows show the deformed shapes $\mathcal{M}_{i\rightarrow i+1}$ and $\mathcal{M}_{i\rightarrow i-1}$ respectively. Using the inter-frame deformations $\mathcal{M}_{i\rightarrow i+1}$, we extract the tracking using a nearest neighbour search as shown in the bottom row. Note that we make no hypothesis on the speed of the motion.
@@ 9.3 Training Details
To allow for more stable learning, the network optimises for $l^{network}$ in gradual steps. First, the network only optimises for $l^{coarse}$ to obtain coarse reconstructions. After $N_{1}$ epochs, assuming that we have roughly located the coarse surface, the refinement step is activated and the network optimises for both $l^{coarse}$ and $l^{fine}$ i.e. for $l^{fusion}$. After $N_{2}$ epochs, given that we have converged to good initial refined surfaces, the association search is activated; the network optimises for both $l^{fusion}$ and $l^{associate}$. Finally, after $N_{3}$ epochs, the deformation search i.e. $l^{deform}$ is activated and the network optimises for $l^{network}$ until convergence. Tab. 9 and Tab. 10 detail this in terms of loss weights for the model trained on D-FAUST [8] and the one trained on DT4D-A [33] respectively.
@@
We train the network with the Adam [28] optimizer and use gradient clipping. We use a learning rate of $10^{-3}$ during the first training epoch, $5\times 10^{-4}$ between the $2^{nd}$ and the $60^{th}$ epoch, $2.5\times 10^{-4}$ between the $60^{th}$ epoch and the $100^{th}$ epoch and $1.25\times 10^{-4}$ after the $100^{th}$ epoch.